OpenStudy (anonymous):

-7z - 7 is less than or the same as - 5z + 12 Solve for z

OpenStudy (anonymous):

\[-7z -7 \le -5z+12\]

OpenStudy (anonymous):

is this ur question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait i got that mixed up let me write it again...

OpenStudy (anonymous):

-7z - 7 \[\le\] -5z + 5

OpenStudy (anonymous):

-2z ≤ 19 \[2z \ge -19\] \[z \ge -\frac{19}{2}\]

OpenStudy (anonymous):

is it ok

OpenStudy (anonymous):

NO!... :/ but thanks for trying..

OpenStudy (anonymous):

how come its not correct?????????

OpenStudy (anonymous):

idk the answer is supposed to be z \[\le\] - ? only one number not two... :/

OpenStudy (anonymous):

it can only be a single digit.. :/

OpenStudy (anonymous):

errr greater than or equal to not less than

OpenStudy (anonymous):

oh i took -5z+12

OpenStudy (anonymous):

\[-7z-7 \le -5z + 5\] \[-7z+5z\le 7 + 5\] \[-2z\le 12\] \[2z\ge -12\] \[z\ge 6\]

OpenStudy (anonymous):

THANKS!!!!!!!!!!!!!! :D!! but i think its - 6

OpenStudy (anonymous):

yeah i again forgot to put -6

OpenStudy (anonymous):

Thanks you are very much appreciated!. ;D